Emily scored 50% higher on her SAT than Joshua did. This means that
A. Joshuas score was 50% lower than Emily's
B. Joshua's score was half of Emily's
C. Joshuas score was 2/3 of Emily's.
I know it isn't A. but i can't figure out how to get the answer to correctly answer it.

Answer by checkley77(12844)

Let Josua's score be 50
Let Emily's score be 50+(.5*50)=7
50/75=2/3 ans. C).